# Beyond Subspace Isolation: Many-to-Many Transformer for Light Field Image Super-resolution

### v1

@@ III-A Problem Formulation
In formal terms, the procedure of LFSR is to upsample the spatial resolution from a low-resolution (LR) LF image $I_{LR}$ to a super-resolved (SR) LF image $I_{SR}$, which serves as an approximation to the corresponding high-resolution (HR) LF image $I_{HR}$. It can be denoted as
@@
where $(U,V)$ and $(W,H)$ stand for the LR image’s angular and spatial resolutions respectively, $C$ denotes the channel dimension, $(u,v)$ indicates an angular location, $(x,y)$ indicates a spatial location, while $r$ represents the scale factor.
A LF image or tensor, as shown in Figure 2(a), can be reshaped into various forms to reveal distinctive subspaces. These encompass a spatial tensor $I_{S}$, revealing the spatial subspace $(W,H)$ depicted in Figure 2(b), an angular tensor $I_{A}$ in the angular subspace $(U,V)$ depicted in Figure 2(c), and EPI tensors $I_{EPI-i}$ which expose the EPI subspace consisting of a spatial and an angular dimension. Figure 2(d) illustrates $(U,W)$, a variant of EPI tensor.
@@ III-B Network Architecture
The architecture of our proposed M2MT-Net is depicted in Figure 3(a). It adopts a streamlined yet effective design, comprising three phases. The initial phase involves preliminary feature extraction, accomplished through $n_{1}$ spatial convolutions. The crux of our architecture, the second component, encompasses a sequence of $n_{2}$ correlation blocks. Each block incorporates two distinctive Transformers, namely a Many-to-Many Transformer (M2MT) and an angular Transformer, operating consecutively. A simplified visualization of a correlation block is given in Figure 3(d). The last phase is pixel generation, which upsamples the extracted features by expanding the channel dimensions by $r^{2}$ times with a $1\times 1$ convolution, followed by a pixel shuffler to increase the spatial resolution from $U\times V\times W\times H\times(C\times r^{2})$ to $U\times V\times rW\times rH\times C$, and lastly, a $3\times 3$ convolution to squeeze the channels. Additionally, residual learning is enforced to allow the network to effectively capture residual information by learning from the differences between the HR and the bicubic-interpolated LR input. Also, within each correlation block, a residual skip connection is utilized to improve the information flow.
@@ III-C Many-to-Many Transformer
A simplified illustration of M2MT is depicted in Figure 3 (b), and a detailed one in Figure 2 (e). Specifically, given a 4D LF tensor in a batch $I\in\mathbb{R}^{B\times U\times V\times W\times H\times C}$, it initiates by reshaping it into a spatial tensor $I_{\tilde{S}}$ in a special form. Diverging from the conventional approach of merging angular dimensions with the batch dimension as shown in Figure 2(b), which yields a spatial tensor $I_{S}$ of dimensions $(BUV\times WH\times C)$ with $WH$ as tokens and $C$ as embedding, we instead merge the angular dimensions with the channel dimensions, resulting in a shape of $(B\times WH\times UVC)$. This approach prepares the tensor for the following correlation encoding process, transforming it into a correlation tensor denoted as $I_{Cor}$ using a linear layer:
$I_{Cor}=L_{encode}(I_{\tilde{S}}),\quad I_{Cor}\in\mathbb{R}^{B\times WH\times C_{Cor}}$ (4)
where $L_{encode}$ represents a linear layer with a projection matrix $W_{encode}\in\mathbb{R}^{UVC\times C_{Cor}}$. The resultant $I_{Cor}$ aggregates the angular correlations at each spatial location. This schema facilitates the succeeding Transformer to invoke the self-attention mechanism in the spatial subspace while concurrently tapping into correlation information from all SAIs. The self-attention mechanism is formally defined as
@@
$=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right),$
@@
In this context, $\hat{I}_{Cor}$ signifies the tensor enhanced by self-attention. $L_{Q}$, $L_{K}$, and $L_{V}$ are the linear layers for calculating queries, keys, and values respectively. Notably, we replace commonly used predefined positional encodings with two $3\times 3$ spatial convolutions to capture locality as suggested by [58].
As last, a linear layer is employed to decode $\hat{I}_{Cor}$ and recover the angular dimensions, resulting in the output feature tensor $\hat{I}$:
$\hat{I}=L_{decode}(\hat{I}_{Cor})$ (6)
@@
$\hat{I}_{Cor}(x,y)\simeq\{I(u,v,\bar{x},\bar{y})\}_{(\bar{x},\bar{y})\in\mathbb{R}^{W\times H}}\$ (8)
Consequently, M2MT is enabled to access the entirety of LF data in a non-local context spatially and angularly with no information remain isolated within the batch dimension:
@@
TABLE I: Quantitative comparisons with the state-of-art methods at 2×2$\times$2 × and 4×4$\times$4 × scales across various datasets. PSNR / SSIM are used as evaluation metrics. The best and second-best results are in bold and underlined.
@@
DistgSSR [22] 2×2$\times$2 × 34.81/0.9787 37.96/0.9796 44.94/0.9949 36.59/0.9859 40.40/0.9942
LFT [31] 2×2$\times$2 × 34.80/0.9781 37.84/0.9791 44.52/0.9945 36.60/0.9855 40.51/0.9941
EPIT [32] 2×2$\times$2 × 34.83/0.9775 38.23/0.9810 45.06/0.9949 36.67/0.9853 42.17/0.9957
M2MT-Net (Ours) 2×2$\times$2 × 35.36/0.9804 38.29/0.9806 45.14/0.9951 36.99/0.9867 40.80/0.9947
M2MT-Net* (Ours) 2×2$\times$2 × 35.56/0.9812 38.48/0.9811 45.39/0.9954 37.19/0.9870 41.23/0.9951
@@
DisgSSR [22] 4×4$\times$4 × 28.99/0.9195 31.38/0.9217 37.56/0.9732 30.99/0.9519 31.65/0.9535
LFT [31] 4×4$\times$4 × 29.26/0.9210 31.46/0.9218 37.63/0.9735 31.21/0.9524 31.86/0.9548
EPIT [32] 4×4$\times$4 × 29.34/0.9197 31.51/0.9231 37.68/0.9737 31.37/0.9526 32.18/0.9571
M2MT-Net (Ours) 4×4$\times$4 × 29.80/0.9277 31.70/0.9252 37.99/0.9750 31.73/0.9556 32.02/0.9560
M2MT-Net* (Ours) 4×4$\times$4 × 29.94/0.9298 31.90/0.9272 38.20/0.9750 31.89/0.9570 32.34/0.9587
@@
HR IINet DistgSSR LFT EPIT M2MT-Net
@@
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
(a) Perforated_Metal_3 20.95/0.7054 20.64/0.7044 21.13/0.7190 18.76/0.5257 22.71/0.8180
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DI = 9.8521 DI = 9.6033 DI = 7.7526 DI = 5.9518 DI = 27.2578
@@
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
(b) Palais_du_Luxembourg 18.86/0.5653 18.62/0.5367 18.90/0.5443 18.96/0.5584 19.68/0.6739
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DI = 8.0609 DI = 7.8105 DI = 4.1552 DI = 6.5223 DI = 21.0431
@@
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
(c) Bicycle 27.01/0.8333 26.85/0.8308 27.14/0.8283 27.21/0.8367 27.37/0.8390
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DI = 11.2056 DI = 6.4649 DI = 4.3732 DI = 6.5302 DI = 19.2823
@@
Figure 5: Visualization of SAI-wise PSNR to demonstrate the distribution of 4×4$\times$4 × LFSR performance. Samples are the same with Figure 4.
III-D Angular Transformer
With M2MT operating in the spatial subspace, an angular transformer is utilized to refine the correlations in the angular subspace. An illustration is depicted in Figure 3(c). This transformer is fundamentally vanilla as in [31, 30], aligning closely with Equation 5, but specifically operates on angular tensors $I_{A}\in\mathbb{R}^{BUV\times WH\times C}$ like Figure 2(c).
Notably, although the M2MT and angular Transformer operate in distinct subspaces, their primary objective converges on the establishment of a comprehensive spatial-angular representation of LF images. In subsequent experiments, we will demonstrate the synergistic benefits of their collaboration, highlighting their complementary nature.
IV Experiments
@@ IV-A Implementation Details
In our experiments, we adhere to the protocols outlined in the commonly used BasicLFSR framework [59] for implementing and benchmarking M2MT-Net. Five public datasets are used, namely EPFL [60], HCInew [61], HCIold [62], INRIA [63], and STFgantry [64]. These datasets contain 70/20/10/35/9 samples for training and 10/4/2/5/2 samples for testing. Following the protocols, we only use the central $5\times 5$ SAIs [59]. For training, each SAI was partitioned into $64\times 64$ or $128\times 128$ patches to serve as HR patches, and $1/2$ or $1/4$ bicubic down-sampling is applied to produce the corresponding LR patches for 2×2$\times$2 × or 4×4$\times$4 × scales, respectively. We use Adam optimizer with a learning rate of $2\times 10^{-4}$ and batches of 4 samples. The training process takes 100 epochs to sufficiently converge. Regarding the hyperparameters, we use $C=48$ across all Transformers and convolutions except Correlation Tensors with $C_{Cor}=128$. The number of initial spatial convolutions is set to 4, and the number of correlation blocks $n_{2}$ is set to 8.
The experiments are conducted on a desktop computer with an Intel i7-11700 4.800GHz 8-core CPU, 32 MB RAM, and an Nvidia GTX 3090 GPU. We implement M2MT-Net using the PyTorch framework. The implementation code and trained models will be released publicly.
@@ IV-B Quantitative Comparisons
A quantitative comparison is conducted to compare M2MT-Net with eight state-of-the-art LFSR methods on the five aforementioned datasets at 2×2$\times$2 × and 4×4$\times$4 × scales. The compared methods include convolution-based LFSSR [21], LF-ATO [48], LF-InterNet [49], LF-IINet [50], DKNet [53], DistgSSR [22], and Transformer-based methods DPT [30], LFT [31], and EPIT [32]. The outcomes are presented in Table I.
@@
It is evident that our M2MT-Net holds a superior position. At both the 2×2$\times$2 × and 4×4$\times$4 × scales, it achieves the highest PSNR across datasets. Notably, on the EPFL dataset, which contains the most testing samples, M2MT-Net surpasses the second-best method, EPIT, by a significant 0.53 dB PSNR gain at the 2×2$\times$2 × scale and 0.46 dB at the 4×4$\times$4 × scale. M2MT-Net ranks second to EPIT in only one dataset, namely STFgantry. This particular outcome can be attributed to the dataset’s distinctive characteristic of exhibiting high disparities, a feature effectively managed by EPIT due to its utilization of the EPI mechanism.
A discernible trend is observed regarding the performance of Transformer-based methods on the EPFL dataset, such as LFT and EPIT, in comparison to convolutional methods, with DistgSSR as the representative. At the 2×2$\times$2 × scale, the superiority of EPIT over DistgSSR is marginal, by just 0.02 dB, with LFT trailing further. Yet, at the 4×4$\times$4 × scale, both LFT and EPIT remarkably outdo DistgSSR by margins of 0.35 dB and 0.27 dB, respectively. This discrepancy underscores the inherent strengths of convolutions in extracting high-frequency information, particularly visual textures, which become advantageous at the 2×2$\times$2 × scale where most lost visual textures are still recoverable. Conversely, Transformers excel in modeling distant pixel or SAI dependencies, rendering them more suitable for the 4×4$\times$4 × scale where a model has to utilize existing spatial and angular cues more effectively. Different from LFT and EPIT, M2MT-Net, with the subspace isolation mitigated, integrates the advantages of both methodologies, thereby achieving significant lead margins of 0.55 dB and 0.81 dB over DistgSSR on the EPFL dataset at both scales.
@@
We also incorporate the geometric self-ensemble strategy, which was initially proposed for single image super-resolution [65], into M2MT-Net to enhance the model performance without introducing additional parameters. The variant is labeled as M2MT-Net* in Table I. Similar to its application in 2D single images, during inference, the strategy transforms the 2D LR by flipping and rotating to construct an ensemble ${T_{i}\overline{I}_{LR}}$, where $T_{i}$ represents a transform function. The SR is generated by executing the network on each member in the ensemble individually, followed by the corresponding inverse transform, and finally averaging the output. The strategy is expressed as
$I_{SR}=\frac{1}{n}\sum_{i=1}^{n}T_{i}^{-1}\mathcal{F}(T_{i}(I_{LR}))$ (10)
@@
where $n$ is the number of transforms. Given that these transforms operate on the reshaped 2D SAI-view LF $\overline{I}_{LR}\in\mathbb{R}^{UV\times WH\times C}$, they take effect on both the spatial and angular spaces synergistically, ensuring the LF structure is not distorted but preserved after transform. The result in Table I demonstrates the advantageous impact brought by the strategy with a roughly 0.10 to 0.25 dB increase in PSNR observed across the datasets and a particular 0.43 dB and 0.32 dB increase on the STFgantry dataset at the 2×2$\times$2 × and 4×4$\times$4 × scales respectively. These findings suggest that the geometric self-ensemble strategy is a valuable addition to compensate and enhance LFSR models.

### v2

@@ III-A Problem Formulation
In formal terms, the procedure of LFSR is to enhance the spatial resolution from a low-resolution (LR) LF image $I_{LR}$ to a super-resolved (SR) LF image $I_{SR}$, which serves as an approximation to the corresponding high-resolution (HR) LF image $I_{HR}$. It can be denoted as
@@
where $\mathcal{F}(\cdot)$ is the super-resolution process, $(U\times V)$ and $(W\times H)$ stand for the LR image’s angular and spatial resolutions, respectively, $C$ denotes the channel dimension, $(u,v)$ indicates an angular location, $(x,y)$ indicates a spatial location, and $r$ represents the scale factor.
A LF image or tensor, as shown in Fig. 2(a), can be reshaped into various forms to reveal distinctive subspaces. These encompass a spatial tensor $I_{S}$, revealing the spatial subspace $(W\times H)$ depicted in Fig. 2(b), an angular tensor $I_{A}$ in the angular subspace $(U\times V)$ depicted in Fig. 2(c), and EPI tensors $I_{EPI}$ which expose an EPI subspace, which consists of a spatial dimension and an angular dimension. Fig. 2(d) illustrates the tensor exposing $(U\times W)$ or $(V\times H)$, two typical EPI subspaces.
@@ III-B Network Architecture
The architecture of our proposed M2MT-Net is depicted in Fig. 3(a). It adopts a streamlined yet effective design comprising three phases. The first phase involves initial feature extraction, accomplished through $n_{1}$ spatial convolution layers. The crux of our architecture, the second phase, encompasses a sequence of $n_{2}$ correlation blocks. Each block incorporates two distinctive Transformers, namely a Many-to-Many Transformer (M2MT) and an angular Transformer, operating consecutively. A simplified visualization of a correlation block is given in Fig. 3(d). The last phase is pixel generation, which upsamples the extracted features by expanding the channel dimensions by $r^{2}$ times with a $1\times 1$ convolution, followed by a pixel shuffler to increase the spatial resolution from $U\times V\times W\times H\times(r^{2}C)$ to $U\times V\times rW\times rH\times C$, and lastly, a $3\times 3$ convolution to squeeze the channels. Additionally, residual learning is enforced to allow the network to effectively capture residual information by learning from the differences between the HR and the bicubic-interpolated LR input. Also, within each correlation block, a residual skip connection is utilized to improve the information flow.
@@ III-C Many-to-Many Transformer
A general Transformer [67] processes an input tensor $X\in\mathbb{R}^{B\times L\times D}$, where $B$ represents the batch dimension, $L$ is a sequence of tokens, and each token is a $D$-dimensional embedding. The Transformer’s self-attention mechanism captures long-range dependencies by integrating information across all $L$ tokens globally.
To handle a 4D LF image $I\in\mathbb{R}^{U\times V\times W\times H\times C}$ using the spatial subspace as tokens, as illustrated in Fig. 2(b), conventional approaches [23, 40] merge the angular subspace with the batch dimension, resulting in a spatial tensor $I_{S}\in\mathbb{R}^{BUV\times WH\times C}$, where $WH$ serves as tokens and $C$ serves as embeddings ($L=WH$ and $D=C$). However, this method leads to subspace isolation, as discussed in Section I.
To address this issue, the proposed M2MT diverges from this conventional approach. A simplified illustration of M2MT is depicted in Fig. 3(b), and a detailed one in Fig. 2(e). Specifically, it initiates by merging the angular subspace with the channel dimensions, yielding a spatial tensor in a special form $I_{\tilde{S}}\in\mathbb{R}^{B\times WH\times UVC}$, which prepares the tensor for the following correlation encoding process. The correlation encoding process transforms $I_{\tilde{S}}$ into a correlation tensor $I_{Cor}\in\mathbb{R}^{B\times WH\times C_{Cor}}$:
$I_{Cor}=L_{encode}(I_{\tilde{S}})$ (4)
where $C_{Cor}$ denotes the number of channels of the correlation tensor, and the correlation encoding process $L_{encode}:\mathbb{R}^{B\times WH\times UVC}\mapsto\mathbb{R}^{B\times WH\times C_{Cor}}$ is implemented through a linear layer (or a fully connected layer) with a weight matrix $W_{encode}\in\mathbb{R}^{UVC\times C_{Cor}}$:
$L_{encode}(X)=W_{encode}X.$ (5)
The resultant $I_{Cor}$ aggregates the angular correlation information at each spatial location into a compact feature representation at a reduced dimensionality of $C_{Cor}$. This schema facilitates the succeeding Transformer to invoke the self-attention mechanism in the spatial subspace while concurrently tapping into the correlation information from all SAIs ($L=WH$ and $D=C_{Cor}$). The self-attention mechanism is formally defined as
@@
$=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{D}}\right)V,$
@@
In this context, $\hat{I}_{Cor}$ signifies the tensor enhanced by self-attention. $L_{Q}$, $L_{K}$, and $L_{V}$ are the linear layers for calculating queries, keys and values ($Q$, $K$ and $V$), respectively, and $D$ is their channel number. Notably, we replace commonly used predefined positional encodings with two $3\times 3$ spatial convolutions to capture locality as suggested by [68].
Finally, $\hat{I}_{Cor}$ undergoes the correlation decoding process to restore the angular subspace. This process mirrors the correlation encoding process in Equation 4 and 5. However, it operates in reverse, using a linear layer $L_{decode}:\mathbb{R}^{B\times WH\times C_{Cor}}\mapsto\mathbb{R}^{B\times WH\times UVC}$ with a weight matrix $W_{decode}\in\mathbb{R}^{C_{Cor}\times UVC}$. The output tensor $\hat{I}$ is then generated as follows:
$\hat{I}=L_{decode}(\hat{I}_{Cor}),$ (7)
$L_{decode}(X)=W_{decode}X.$ (8)
@@
$\hat{I}_{Cor}(x,y)\simeq\{I_{Cor}(\bar{x},\bar{y})\}_{(\bar{x},\bar{y})\in\mathbb{R}^{W\times H}}\$ (10)
As a result, M2MT is enabled to access the entirety of LF data in a non-local context spatially and angularly with no information remaining isolated within the batch dimension:
@@
III-D Angular Transformer
While M2MT achieves interactions in the spatial subspace, it remains crucial to engage an angular component to facilitate interactions within the angular subspace. To this end, an angular transformer is utilized to refine the correlation in the angular subspace. An illustration is depicted in Fig. 3(c). This Transformer is fundamentally vanilla as in [40, 39], aligning closely with Equation 6, but specifically operates on angular tensors $I_{A}\in\mathbb{R}^{BUV\times WH\times C}$ as depicted in Fig. 2(c). The channel number of key, query, and value is set to $D=C$.
Notably, although the M2MT and angular Transformer operate in distinct subspaces, their primary objective converges on the establishment of a comprehensive spatial-angular representation of LF images. In Section IV-G1 and TABLE III, we demonstrate that M2MT alone establishes a solid foundation of a competitive network, incorporating angular Transformers offers a complementary effect that further enhances M2MT-Net’s overall performance by effectively managing angular interactions.
IV Experiments
TABLE I: Quantitative comparisons with the state-of-art methods at the 2×2$\times$2 × and 4×4$\times$4 × scales across various datasets. PSNR / SSIM are used as evaluation metrics. The best and second-best results are in bold and underlined, respectively.
@@ IV Experiments
LFSSR-SAV [27] 4×4$\times$4 × 34.62/0.9772 37.42/0.9776 44.22/0.9942 36.36/0.9849 38.69/0.9914
DistgSSR [22] 2×2$\times$2 × 34.81/0.9787 37.96/0.9796 44.94/0.9949 36.58/0.9859 40.40/0.9942
LFT [40] 2×2$\times$2 × 34.78/0.9776 37.77/0.9788 44.63/0.9947 36.54/0.9853 40.41/0.9941
EPIT [41] 2×2$\times$2 × 34.85/0.9775 38.23/0.9810 45.08/0.9949 36.68/0.9852 42.17/0.9957
HLFSR [62] 4×4$\times$4 × 35.31/0.9800 38.32/0.9807 44.98/0.9950 37.06/0.9867 40.85/0.9947
LF-DET [23] 4×4$\times$4 × 35.20/0.9794 38.22/0.9803 44.92/0.9949 36.88/0.9862 41.56/0.9953
M2MT-Net (Ours) 2×2$\times$2 × 35.64/0.9815 38.43/0.9810 45.38/0.9953 37.22/0.9870 40.99/0.9949
M2MT-Net* (Ours) 2×2$\times$2 × 35.82/0.9822 38.62/0.9816 45.58/0.9955 37.40/0.9873 41.39/0.9953
@@
LFSSR-SAV [27] 4×4$\times$4 × 29.37/0.9223 31.45/0.9217 37.50/0.9721 31.27/0.9531 31.36/0.9505
DisgSSR [22] 4×4$\times$4 × 28.99/0.9195 31.38/0.9217 37.56/0.9732 30.99/0.9519 31.65/0.9534
LFT [40] 4×4$\times$4 × 29.33/0.9196 31.36/0.9205 37.59/0.9731 31.30/0.9515 31.62/0.9548
EPIT [41] 4×4$\times$4 × 29.31/0.9196 31.51/0.9231 37.68/0.9737 31.35/0.9526 32.18/0.9570
HLFSR [62] 4×4$\times$4 × 29.20/0.9222 31.57/0.9238 37.78/0.9742 31.24/0.9534 31.64/0.9537
LF-DET [23] 4×4$\times$4 × 29.42/0.9220 31.51/0.9227 37.76/0.9739 31.34/0.9528 32.02/0.9561
M2MT-Net (Ours) 4×4$\times$4 × 29.85/0.9284 31.76/0.9264 37.98/0.9749 31.77/0.9563 32.20/0.9584
M2MT-Net* (Ours) 4×4$\times$4 × 29.96/0.9300 31.94/0.9279 38.21/0.9758 31.87/0.9572 32.45/0.9602
@@
HR DistgSSR HLFSR LFT EPIT LF-DET M2MT-Net
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
(a) Perforated_Metal_3 24.14/0.7053 24.57/0.7216 24.24/0.7040 24.18/0.6989 24.81/0.7218 25.49/0.7573
@@
DI = 8.7671 DI = 12.5190 DI = 5.8032 DI = 8.3895 DI = 23.3273 DI = 25.1975
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
(b) Palais_du_Luxembourg 22.69/0.7237 22.75/0.7295 22.70/0.7218 22.58/0.7115 22.87/0.7325 23.16/0.7466
@@
DI = 8.0693 DI = 12.5424 DI = 5.1459 DI = 8.2841 DI = 23.7106 DI = 25.2688
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
(c) Bicycle 26.32/0.6822 26.54/0.6913 26.33/0.6806 26.46/0.6858 26.55/0.6885 26.85/0.7001
@@
DI = 9.1141 DI = 13.1903 DI = 6.6828 DI = 8.2787 DI = 18.7618 DI = 19.7059
@@
Figure 5: Visualization of SAI-wise PSNR to demonstrate the distribution of 4×4$\times$4 × LFSR performance. The compared samples are the same with Fig. 4.
@@ IV-A Implementation Details
In our experiments, M2MT-Net is implemented using the deep learning framework PyTorch [69]. We adhere to the protocols outlined in the widely used BasicLFSR framework [70] to conduct evaluations in a fair and consistent manner. Five public datasets are used, namely EPFL [71], HCInew [72], HCIold [73], INRIA [74], and STFgantry [75]. These datasets contain 70/20/10/35/9 samples for training and 10/4/2/5/2 samples for testing. Following the protocols, we only use the central $5\times 5$ SAIs [70]. For training, each SAI was partitioned into $64\times 64$ or $128\times 128$ patches to serve as HR patches, and $1/2$ or $1/4$ bicubic down-sampling is applied to produce the corresponding LR patches for 2×2$\times$2 × or 4×4$\times$4 × scales, respectively. We use Adam optimizer with a learning rate of $2\times 10^{-4}$ and batches of 4 samples. The training process takes 60 epochs to converge and five epochs to fine-tune. Regarding the hyperparameters, empirically, we use $C=48$ across all Transformers and convolutions except M2MT’s correlation tensors and query, key and values with $C_{Cor}=D=128$. The number of spatial convolutions in the initial feature extraction $n_{1}$ is set to 4. The number of correlation blocks $n_{2}$ is set to 9 for the 2×2$\times$2 × scale and 8 for the 4×4$\times$4 × scale.
The experiments are conducted on a computer with an Intel i7-11700 4.800GHz 8-core CPU, 32 MB RAM, and an Nvidia GTX 3090 GPU. The implementation code and trained models are released publicly at https://huzexi.github.io/.
@@ IV-B Quantitative Comparisons
A quantitative comparison is conducted to compare M2MT-Net with eight state-of-the-art LFSR methods on the five aforementioned datasets at the 2×2$\times$2 × and 4×4$\times$4 × scales. The compared methods include convolution-based LFSSR [21], LF-ATO [56], LF-InterNet [57], LF-IINet [58], DKNet [61], LFSSR-SAV [27], DistgSSR [22], HLFSR [62] and Transformer-based DPT [39], LFT [40], EPIT [41] and LF-DET [23]. Their publicly released weight files are utilized to conduct this comparison. The outcomes are presented in Table I.
@@
It is evident that our M2MT-Net holds a superior position. At both the 2×2$\times$2 × and 4×4$\times$4 × scales, it achieves the highest PSNR across almost all datasets. Notably, on the EPFL dataset, which contains the most testing samples, M2MT-Net surpasses the second-best method, LF-DET, by a significant 0.43 dB PSNR gain at the 4×4$\times$4 × scale and 0.33 dB at the 2×2$\times$2 × scale. M2MT-Net ranks third on only one dataset, STFgantry, at the $2\times scale$. This particular outcome can be attributed to the dataset’s distinctive characteristic of exhibiting high disparities, which is effectively addressed by the EPI mechanism of EPIT and HLFSR and LF-DET’s multi-scale angular modeling. However, their advantage does not extend to the 4×4$\times$4 × scale, where M2MT-Net reclaims its lead, surpassing EPIT, HLFSR and LF-DET by margins of 0.02 dB, 0.56 dB and 0.18 dB, respectively.
A notable trend is observed regarding the performance between the 2×2$\times$2 × and 4×4$\times$4 × scales. While the PSNR advantage of M2MT-Net over the second best methods at 2×2$\times$2 × scale is relatively modest from 0.11 to 0.30 dB, the gap widens significantly at the 4×4$\times$4 × scale, ranging from 0.19 to 0.43 dB. This discrepancy highlights the inherent strengths of M2MT-Net in handling the more challenging 4×4$\times$4 × scale, where more details are lost due to down-sampling, requiring the model to utilize existing spatial and angular cues more effectively.
@@
We also incorporate the geometric self-ensemble strategy, which was initially proposed for single image super-resolution [76], into M2MT-Net to enhance the model performance without introducing additional parameters. The variant is labeled as M2MT-Net* in Table I. Similar to its application in 2D single images, during inference, the strategy transforms the 2D LR by flipping and rotating to construct an ensemble $\{T_{i}(\overline{I}_{LR})\}$, where $T_{i}$ represents a transform function. The SR is generated by executing the network on each member in the ensemble individually, followed by the corresponding inverse transform, and finally, averaging the output. The strategy is expressed as
$I_{SR}=\frac{1}{n}\sum_{i=1}^{n}T_{i}^{-1}(\mathcal{F}(T_{i}(I_{LR})))$ (12)
@@
where $n$ is the number of transforms. The transforms take place on the spatial and angular subspaces synergistically to ensure that the LF structure is not distorted but preserved after the transforms. The result in Table I demonstrates the advantageous impact brought by the strategy with a roughly 0.10 to 0.25 dB increase in PSNR observed across the datasets at both scales and a particular 0.40 dB increase on the STFgantry dataset at the 2×2$\times$2 × scales respectively. These findings suggest that the geometric self-ensemble strategy is a valuable addition to compensate for LFSR models.